# Lennart Berg and Torben Wiedemann

### v1

@@ 6 The parametrisation theorem
The general assertion can be deduced from the special case that $\sigma_{\bar{\beta}}$ and $\sigma_{\bar{\gamma}}$ stabilise $\alpha$ and that $\bar{\gamma}$ is the empty word. Then there exist $\rho_{1},\ldots,\rho_{n}\in H_{3}\cap\alpha^{\perp}$ with $\sigma_{\bar{\beta}}=\sigma_{\rho_{1}}\cdots\sigma_{\rho_{n}}$. Using induction, we can reduce to the case that $n=1$, or in other words, that $\sigma_{\bar{\beta}}=\sigma_{\rho}$ for a root $\rho\in\alpha^{\perp}$. Using square-compatibility, the braid-invariance of $\eta$, the braid relations for $(w_{\delta})_{\delta\in\Delta}$ and the solution of the word problem in Coxeter groups, we can replace $\bar{\beta}$ by any word $\bar{\beta}^{\prime}$ with $\sigma_{\bar{\beta}}=\sigma_{\bar{\beta}^{\prime}}$. In particular, we can replace $\bar{\beta}$ by a $\Delta$-expression $\bar{\beta}^{\prime}$ of $\rho$. Then $w_{\bar{\beta}^{\prime}}$ centralises $U_{\alpha}$ by 6.2 and $\eta(\alpha,\bar{\beta}^{\prime})=1_{A}$ because $\eta$ is adjacency-trivial. This finishes the proof. ∎
The following result is precisely the content of the parametrisation theorem [Wie23, 4.5.16] for the special case $\Phi=H_{3}$.
Theorem 6.15 ([Wie23, 4.5.16]).
Under the assumptions of 6.14, there exist an abelian group $(\mathcal{S},+)$, an action $\omega\colon A\times\mathcal{S}\rightarrow\mathcal{S},(b,r)\mapsto b.r$ of $A$ on the underlying set of $\mathcal{S}$ and a parametrisation $(\theta_{\alpha}\colon\mathcal{S}\rightarrow U_{\alpha})_{\alpha\in H_{3}}$ of $G$ by $(A,\mathcal{S})$ with respect to $\eta$ and $(w_{\delta})_{\delta\in\Delta}$ such that the twisting actions of $A$ on the root groups are induced by $\omega$. That is, we have $b.\theta_{\alpha}(s)=\theta_{\alpha}(b.s)$ for all $\alpha\in H_{3}$, $s\in\mathcal{S}$ and $b\in A$.
Sketch of the proof.
@@
Fix a root $\alpha_{0}\in H_{3}$. Choose a group $\mathcal{S}$ which is isomorphic to $U_{\alpha_{0}}$ and an isomorphism $\theta_{\alpha_{0}}\colon\mathcal{S}\rightarrow U_{\alpha_{0}}$. We define an action of $A$ on $\mathcal{S}$ by $b.s\coloneqq\theta_{\alpha_{0}}^{-1}(b.\theta_{\alpha_{0}}(s))$ for all $b\in A$ and $s\in\mathcal{S}$. For any root $\beta$, we define $\theta_{\beta}\colon\mathcal{S}\rightarrow U_{\beta},s\mapsto\theta_{\alpha_{0}}(s)^{w_{\bar{\gamma}}}$ where $\bar{\gamma}$ is any word over Δ($\Delta\cup(-\Delta)$ with $\alpha_{0}^{\sigma(\bar{\gamma})}=\beta$ and $\eta(\alpha,\bar{\gamma})=1_{A}$. Such a word $\bar{\gamma}$ exists by the completeness of $\eta$, and $\theta_{\beta}$ does not depend on the choice of $\bar{\gamma}$ by 6.14. One can verify that this construction has the desired properties. ∎
In order to apply the parametrisation theorem 6.15 to $(G,(U_{\alpha})_{\alpha\in H_{3}})$, we need to find an appropriate twisting group $A$ and a $\Delta$-parity map $\eta$. The former was constructed in 6.10. The latter can be “read off” from the $H_{3}$-graded group in Section 4.

### v2

@@ 6 The parametrisation theorem
Fix a root $\alpha_{0}\in H_{3}$. Choose a group $\mathcal{S}$ which is isomorphic to $U_{\alpha_{0}}$ and an isomorphism $\theta_{\alpha_{0}}\colon\mathcal{S}\rightarrow U_{\alpha_{0}}$. We define an action of $A$ on $\mathcal{S}$ by $b.s\coloneqq\theta_{\alpha_{0}}^{-1}(b.\theta_{\alpha_{0}}(s))$ for all $b\in A$ and $s\in\mathcal{S}$. For any root $\beta$, we define $\theta_{\beta}\colon\mathcal{S}\rightarrow U_{\beta},s\mapsto\theta_{\alpha_{0}}(s)^{w_{\bar{\gamma}}}$ where $\bar{\gamma}$ is any word over Δ($\Delta\cup(-\Delta)$ with $\alpha_{0}^{\sigma(\bar{\gamma})}=\beta$ and $\eta(\alpha,\bar{\gamma})=1_{A}$. Such a word $\bar{\gamma}$ exists by the completeness of $\eta$, and $\theta_{\beta}$ does not depend on the choice of $\bar{\gamma}$ by 6.21. One can verify that this construction has the desired properties. ∎
Notation 6.23.
For the rest of this section, we assume that $\Phi=H_{\ell}$ for some $\ell\in\{3,4\}$, that $X$ is defined as in 4.1 and that $\Delta$ is in standard order. We write $\Delta=(\rho_{1},\rho_{2},\rho_{3})$ if $\ell=3$ and $\Delta=(\rho_{0},\rho_{1},\rho_{2},\rho_{3})$ if $\ell=4$.
Conclusion 6.24.
Denote by $A\coloneqq\{\pm 1\}^{2}$ the standard twisting group for $(G,(w_{\delta})_{\delta\in\Delta})$ from 6.14 and by $\eta\colon H_{\ell}\rightarrow A$ the standard parity map from 6.7. By 6.16, $(G,(U_{\alpha})_{\alpha\in H_{\ell}})$ and $\eta$ satisfy the assumptions in 6.22. Hence there exist an abelian group $(\mathcal{S},+)$ on which $A$ acts and a parametrisation $(\theta_{\alpha})_{\alpha\in H_{\ell}}$ of $G$ by $(A,\mathcal{S})$ with respect to $\eta$ and $(w_{\delta})_{\delta\in\Delta}$. Putting $s^{*}\coloneqq(-1,1).s$ for all $s\in\mathcal{S}$, we can summarise the property of being a parametrisation as follows:
for all $\alpha\in H_{\ell}$, $\delta\in\Delta$ and $s\in\mathcal{S}$. Note that $(s^{*})^{*}=s$ for all $s\in\mathcal{S}$ because $(-1,1)^{2}=(1,1)$. Further, we define a multiplication on $\mathcal{S}$ by
$a\cdot b\coloneqq\theta_{\rho_{1}+\rho_{2}}^{-1}([\theta_{\rho_{1}}(a),\theta_{\rho_{2}}(b)]).$ (6.26)
It satisfies the distributive laws by 3.9 because $(\rho_{1},\rho_{2})$ is an $A_{2}$-pair. Finally, we define $1_{\mathcal{S}}$ to be the element of $\mathcal{S}$ which satisfies $w_{\rho_{2}}\in U_{-\rho_{2}}\theta_{\rho_{2}}(1_{\mathcal{S}})U_{-\rho_{2}}$, which is uniquely determined by 5.3 (iii). By 5.3 (iv), we have
for all $s\in\mathcal{S}$. A similar computation (see [Wie24, 5.7.12]) shows that $1_{\mathcal{S}}\cdot r=r$ for all $r\in\mathcal{S}$ as well, so that $(\mathcal{S},+,\cdot)$ is a nonassociative ring with unit $1_{\mathcal{S}}$.
Goal 6.27.
We will show in Section 9 that the multiplication on $\mathcal{S}$ from 6.24 is associative (9.26) and commutative (9.3), that $\mathcal{S}$ decomposes as $\mathcal{S}=\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ for a pair of isomorphic ideals $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ of $\mathcal{S}$ (9.21, 9.24) and that $(r,s)^{*}=(-r,s)$ for all $(r,s)\in\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ (9.27). Further, we will show that the commutator relations in $G$ are precisely those given in Figure 5 (9.29, 9.30). This says precisely that the parametrisation in 6.24 is a standard coordinatisation in the following sense.
Definition 6.28.
Let $\mathcal{R}$ be a commutative ring. A standard coordinatisation of $(G,(U_{\alpha})_{\alpha\in H_{\ell}})$ by $\mathcal{R}\times\mathcal{R}$ (or, in the terminology of [Wie24, 4.7.5], a coordinatisation with standard signs) is a family $(\theta_{\alpha}\colon(\mathcal{R}\times\mathcal{R},+)\rightarrow U_{\alpha})_{\alpha\in H_{\ell}}$ of root isomorphisms with the following properties.
  1. (i)
    $w_{\delta}^{\prime}\coloneqq\theta_{-\delta}(-1_{\mathcal{R}},-1_{\mathcal{R}})\theta_{\delta}(1_{\mathcal{R}},1_{\mathcal{R}})\theta_{-\delta}(-1_{\mathcal{R}},-1_{\mathcal{R}})$ is a $\delta$-Weyl element for all $\delta\in\Delta$.
  2. (ii)
    $(\theta_{\alpha})_{\alpha\in H_{\ell}}$ is a parametrisation of $(G,(U_{\alpha})_{\alpha\in H_{\ell}})$ (in the sense of 6.11) by the standard parameter system (×,{±$(\mathcal{R}\times\mathcal{R},\{\pm 1\}\times\{\pm 1\})$ from 6.12 with respect to the standard parity map from 6.7 and $(w_{\delta}^{\prime})_{\delta\in\Delta}$.
  3. (iii)
    The same commutator relations as in 4.29 are satisfied.
We also say that $(\theta_{\alpha})_{\alpha\in H_{\ell}}$ is a standard coordinatisation with respect to $(w_{\delta}^{\prime})_{\delta\in\Delta}$.
Example 6.29.
The families $(\theta_{\alpha}^{H})_{\alpha\in H_{3}}$ and $(\theta_{\alpha}^{H})_{\alpha\in H_{\ell}}$ from 4.26 are standard coordinatisations of $\operatorname{Ch}(D_{6},\mathcal{R})$ and $\operatorname{Ch}(E_{8},\mathcal{R})$, respectively. Here $\mathcal{R}$ is any commutative ring.
7 Standard coordinatisations and the Steinberg presentation
Notation 7.1.
In this section, we fix $\ell\in\{3,4\}$ and a commutative ring $\mathcal{R}$. We put $X\coloneqq D_{6}$ if $\ell=3$ and $X\coloneqq E_{8}$ if $\ell=4$.
The purpose of this section is to collect some properties of $H_{\ell}$-graded groups with a standard coordinatisation. They are not needed for the proof of our main result 9.33, by which every $H_{\ell}$-graded group has a standard coordinatisation.
The following remark is analogous to [Wie24, 10.4.16], which is about $F_{4}$-gradings.
Remark 7.2.
Let $(G,(U_{\alpha})_{\alpha\in H_{\ell}})$ be an $H_{\ell}$-graded group with a standard coordinatisation $(\theta_{\alpha})_{\alpha\in H_{\ell}}$ by $\mathcal{R}\times\mathcal{R}$. Recall from 3.17 that for all non-proportional $\alpha,\gamma\in H_{\ell}$ and all $\beta\in\mathord{]\alpha,\gamma[}$, we have a commutation map $\psi_{\alpha,\gamma}^{\beta}\colon M\times M\rightarrow M$ such that
$[\theta_{\alpha}(r),\theta_{\gamma}(s)]=\prod_{\beta\in\mathord{]\alpha,\beta[}}\theta_{\beta}\big{(}\psi_{\alpha,\gamma}^{\beta}(r,s)\big{)}$
for all $r,s\in M$ where the product is taken in the (unique) interval ordering on $\mathord{]\alpha,\gamma[}$ starting from $\alpha$. Axiom 6.28 iii says precisely that the commutation map $\psi_{\alpha,\gamma}^{\beta}$ is prescribed for all $\alpha,\gamma\in E_{2}^{+}(\Delta)$ and $\beta\in\mathord{]\alpha,\gamma[}$, but no explicit statement is made about the case $\alpha,\gamma\notin E_{2}^{+}(\Delta)$. However, let $(\xi,\zeta)$ be an arbitrary pair of non-proportional roots in $H_{\ell}$. Then by 2.9, there exists $u\in\operatorname{Weyl}(H_{\ell})$ such that $\alpha\coloneqq\xi^{u}$ and $\gamma\coloneqq\zeta^{u}$ lie in $E_{2}^{+}(\Delta_{H})$. Choose $\delta_{1},\ldots,\delta_{k}\in\Delta$ such that $u^{-1}=\sigma(\delta_{1})\cdots\sigma(\delta_{k})$ and put $w\coloneqq w_{\delta_{1}}^{\prime}\cdots w_{\delta_{k}}^{\prime}$ where $w_{\delta_{1}}^{\prime},\ldots,w_{\delta_{k}}^{\prime}$ are the Weyl elements from Axiom 6.28 i. Then it follows from the equation above that
$[\theta_{\alpha}(r)^{w},\theta_{\gamma}(s)^{w}]=\prod_{\beta\in\mathord{]\alpha,\beta[}}\theta_{\beta}\big{(}\psi_{\alpha,\gamma}^{\beta}(r,s)\big{)}^{w}$
is the commutator formula for the pair $(U_{\xi},U_{\zeta})$. By Axiom 6.28 ii, it looks just like the commutator formula for $(U_{\alpha},U_{\gamma})$, except that the action of the twisting group may introduce or change some signs. The precise signs that appear in the new formula can be computed using the values of the standard parity map, which are given in Figures 412 and 13. (We will see examples of such computations in 9.30 and 10.2.) It follows that the commutation maps $\psi_{\xi,\zeta}^{\beta}$ with respect to a standard coordinatisation are prescribed for all non-proportional $\xi,\zeta\in H_{\ell}$ and all $\beta\in\mathord{]\xi,\zeta[}$ and that they may be computed explicitly. They are called the standard commutation maps (with respect to $M=\mathcal{R}\times\mathcal{R}$). Note that they are maps $M\times M\rightarrow M$ which depend only on $\mathcal{R}$ but not on $G$.
We now construct a universal $H_{\ell}$-graded group for every commutative ring. Every $H_{\ell}$-graded group is the quotient of one such group by 7.5 and our main result 9.33.
Definition 7.3.
The Steinberg group of type $H_{\ell}$ over $\mathcal{R}$ is the abstract group $\operatorname{St}_{H_{\ell}}(\mathcal{R})$ which is generated by elements $\theta_{\alpha}^{\text{St}}(r,s)$ for $\alpha\in H_{\ell}$ and $r,s\in\mathcal{R}$ with respect to the following relations:
  1. (a)
    $\theta_{\alpha}^{\text{St}}(r+r^{\prime},s+s^{\prime})=\theta_{\alpha}^{\text{St}}(r,s)\theta_{\alpha}^{\text{St}}(r^{\prime},s^{\prime})$ for all $\alpha\in H_{\ell}$ and all $r,r^{\prime},s,s^{\prime}\in\mathcal{R}$.
  2. (b)
    For all non-proportional $\alpha,\gamma\in H_{\ell}$, we require the commutator relation
    $[\theta_{\alpha}^{\text{St}}(r),\theta_{\gamma}^{\text{St}}(s)]=\prod_{\beta\in\mathord{]\alpha,\beta[}}\theta_{\beta}^{\text{St}}\big{(}\psi_{\alpha,\gamma}^{\beta}(r,s)\big{)}$
    for all $r,s\in\mathcal{R}\times\mathcal{R}$ where the product is taken in the (unique) interval ordering on $\mathord{]\alpha,\gamma[}$ starting from $\alpha$ and where
    $\psi_{\alpha,\gamma}^{\beta}\colon(\mathcal{R}\times\mathcal{R})\times(\mathcal{R}\times\mathcal{R})\rightarrow\mathcal{R}\times\mathcal{R}$
    denotes the standard commutation map with respect to $\mathcal{R}\times\mathcal{R}$ (see 7.2).
For all $\alpha\in H_{\ell}$, we define a subgroup $U_{\alpha}^{\text{St}}\coloneqq\{\theta_{\alpha}^{\text{St}}(r)\mid r\in\mathcal{R}\times\mathcal{R}\}$ and we consider $\theta_{\alpha}^{\text{St}}$ as a surjective homomorphism from $\mathcal{R}\times\mathcal{R}$ to $U_{\alpha}^{\text{St}}$.
Remark 7.4.
The Steinberg group is functorial in $\mathcal{R}$: If ϕ::italic-$\phi\colon\mathcal{R}\rightarrow\mathcal{R}^{\prime}$ is a homomorphism of commutative rings, then there is a unique homomorphism $\operatorname{St}(\varphi)\colon\operatorname{St}_{H_{\ell}}(\mathcal{R})\rightarrow\operatorname{St}_{H_{\ell}}(\mathcal{R}^{\prime})$ with $\operatorname{St}(\varphi)\circ\theta_{\alpha,\mathcal{R}}^{\text{St}}=\theta_{\alpha,\mathcal{R}^{\prime}}^{\text{St}}$ for all $\alpha\in H_{\ell}$.
Lemma 7.5.
Let $(G,(U_{\alpha})_{\alpha\in H_{\ell}})$ be an $H_{\ell}$-graded group with a standard coordinatisation $(\theta_{\alpha})_{\alpha\in H_{\ell}}$ by $\mathcal{R}\times\mathcal{R}$. Then there exists a unique surjective homomorphism $\pi\colon\operatorname{St}_{H_{\ell}}(\mathcal{R})\rightarrow G$ such that $\pi(\theta_{\alpha}^{\text{St}}(r,s))=\theta_{\alpha}(r,s)$ for all $\alpha\in H_{\ell}$ and all $r,s\in\mathcal{R}$. Further, for any positive system $\Pi$ in $H_{\ell}$, the restriction $\pi\colon U_{\Pi}^{\text{St}}\rightarrow U_{\Pi}$ is bijective.
Proof.
The existence and uniqueness of $\pi$ follows from the universal property of groups defined by generators and relations. Now let $\Pi$ be a positive system in $H_{\ell}$ and let $\alpha_{1},\ldots,\alpha_{k}$ be an ordering of the roots in $\Pi$ with the properties in 3.12. Let $x\in U_{\Pi}^{\text{St}}$ such that $\pi(x)=1$. Since $\operatorname{St}_{H_{\ell}}(\mathcal{R})$ has $H_{\ell}$-commutator relations by definition, it follows from 3.12 a that there exist $r_{1},\ldots,r_{k}\in\mathcal{R}\times\mathcal{R}$ with $x=\prod_{i=1}^{k}\theta_{\alpha_{i}}^{\text{St}}(r_{i})$. Then $1=\pi(x)=\prod_{i=1}^{k}\theta_{\alpha_{i}}(r_{i})$, which by 3.12 b implies that $r_{i}=0$ for all $i\in\{1,\ldots,k\}$. Hence $x=1$, so $\pi\colon U_{\Pi}^{\text{St}}\rightarrow U_{\Pi}$ is bijective. ∎
Lemma 7.6.
The maps $(\theta_{\alpha}^{\text{St}})_{\alpha\in H_{\ell}}$ are injective and $(\operatorname{St}_{H_{\ell}}(\mathcal{R}),(U_{\alpha}^{\text{St}})_{\alpha\in H_{\ell}})$ satisfies Axiom 3.4 (iv).
Proof.
Choose an $H_{\ell}$-graded group $(G,(U_{\alpha})_{\alpha\in H_{\ell}})$ with a standard coordinatisation $(\theta_{\alpha})_{\alpha\in H_{\ell}}$ by $\mathcal{R}\times\mathcal{R}$. (For example, one of the Chevalley groups from Section 4.) Let $\pi\colon\operatorname{St}_{H_{\ell}}(\mathcal{R})\rightarrow G$ be the homomorphism from 7.5. Let $\Pi$ be a positive system in $H_{\ell}$, let $\alpha\in-\Pi$ and let $x\in U_{\alpha}^{\text{St}}\cap U_{\Pi}^{\text{St}}$. Then $\pi(x)\in U_{\alpha}\cap U_{\Pi}=\{1_{G}\}$. As $\pi$ is injective on $U_{\Pi}^{\text{St}}$ by 7.5, it follows that $x=1$. Hence $\operatorname{St}_{H_{\ell}}(\mathcal{R})$ satisfies Axiom 3.4 (iv). Further, for each $\alpha\in H_{\ell}$, the injectivity of $\theta_{\alpha}$ implies the injectivity of $\theta_{\alpha}^{\text{St}}$ because $\theta_{\alpha}=\pi\circ\theta_{\alpha}^{\text{St}}$. ∎
The existence of Weyl elements in $\operatorname{St}_{H_{\ell}}(\mathcal{R})$ essentially follows from the existence of Weyl elements in some $H_{\ell}$-graded group with a standard coordinatisation.
Lemma 7.7.
Let $\alpha\in H_{\ell}$ and let $r,s\in\mathcal{R}$ be invertible. Then
$w_{\alpha}^{\text{St}}(r,s)\coloneqq\theta_{-\alpha}^{\text{St}}(-r^{-1},-s^{-1})\theta_{\alpha}^{\text{St}}(r,s)\theta_{-\alpha}^{\text{St}}(-r^{-1},-s^{-1})$
is an $\alpha$-Weyl element in $\operatorname{St}_{H_{\ell}}(\mathcal{R})$.
Proof.
Let $\beta\in H_{\ell}$ and $w_{\alpha}\coloneqq w_{\alpha}^{\text{St}}(r,s)$. We have to show that $(U_{\beta}^{\text{St}})^{w_{\alpha}}=U_{\beta^{\sigma(\beta)}}^{\text{St}}$. Since $r,s$ are arbitrary and $w_{\alpha}^{-1}=w_{\alpha}(-r,-s)$, it suffices to show that $(U_{\beta}^{\text{St}})^{w_{\alpha}}\subseteq U_{\beta^{\sigma(\alpha)}}^{\text{St}}$. First, assume that $\beta\notin\{\alpha,-\alpha\}$. Let $x_{\beta}\in U_{\beta}^{\text{St}}$. Let $\Pi$ be the unique positive system in the rank-2 subsystem $\Psi_{\alpha,\beta}$ such that $\alpha,\beta\in\Pi$ and such that the root base of $\Psi_{\alpha,\beta}$ associated to $\Pi$ contains $\alpha$. Put $P\coloneqq\Pi\setminus\{\alpha\}$. Then for all $g\in U_{\alpha}^{\text{St}}\cup U_{-\alpha}^{\text{St}}$ and all $\gamma\in P$, we have $(U_{\gamma}^{\text{St}})^{g}\subseteq[g,U_{\gamma}^{\text{St}}]U_{\gamma}^{\text{St}}\subseteq U_{P}^{\text{St}}$ by 2.2 (i). Hence
$(U_{P}^{\text{St}})^{w_{\alpha}}\subseteq U_{P}^{\text{St}}\subseteq U_{\Pi}^{\text{St}}.$
Now consider the Chevalley group $(\operatorname{Ch}(X,\mathcal{R}),(\tilde{U}_{\alpha}^{H})_{\alpha\in X})$ from Section 4 and the surjective homomorphism $\pi\colon\operatorname{St}_{H_{\ell}}(\mathcal{R})\rightarrow\operatorname{Ch}(X,\mathcal{R})$ from 7.5. Since $\pi(w_{\alpha})$ is an $\alpha$-Weyl element in $\operatorname{Ch}(X,\mathcal{R})$ by 4.27, we have $\pi(x_{\beta}^{w_{\alpha}})\in\tilde{U}_{\beta^{\sigma(\alpha)}}^{H}$. Since the restriction $\pi\colon U_{\Pi}^{\text{St}}\rightarrow\tilde{U}_{\Pi}^{H}$ is an isomorphism and $x_{\beta}^{w_{\alpha}}\in(U_{P}^{\text{St}})^{w_{\alpha}}\subseteq U_{\Pi}^{\text{St}}$, we infer that $x_{\beta}^{w_{\alpha}}\in U_{\beta^{\sigma(\alpha)}}^{\text{St}}$, as desired.
Now assume that $\beta=\alpha$. Choose an $A_{2}$-pair $(\gamma,\gamma^{\prime})$ with $\gamma+\gamma^{\prime}=\alpha$. Recall from 7.2 that up to signs, the commutation map $\psi_{\gamma,\gamma^{\prime}}^{\alpha}$ looks like the map $\psi_{\rho_{0},\rho_{1}}$ in Figure 5. Hence $[U_{\gamma}^{\text{St}},U_{\gamma^{\prime}}^{\text{St}}]=U_{\alpha}^{\text{St}}$. It follows that
$(U_{\alpha}^{\text{St}})^{w_{\alpha}}=[(U_{\gamma}^{\text{St}})^{w_{\alpha}},(U_{\gamma^{\prime}}^{\text{St}})^{w_{\alpha}}]\subseteq[U_{-\gamma^{\prime}}^{\text{St}},U_{-\gamma}^{\text{St}}]=U_{-\alpha}^{\text{St}}.$
The case $\beta=-\alpha$ can be treated similarly. Thus $w_{\alpha}^{\text{St}}(r,s)$ is a Weyl element. ∎